OpenStudy (anonymous):

When the current in one coil of a transformer is increasing by 10amps each second, an induced voltage of 120 V is measured in the other coil. Calculate the mutual inductance.

OpenStudy (anonymous):

V=Mdi/dt=10M---->M=V/10=120/10=12H